# Parameter estimation of gravitational-wave signals with frequency-dependent antenna responses and higher-modes

### v1

@@ Abstract
We implement frequency-dependent antenna responses and develop likelihood classes (standard likelihood, multibanded likelihood, and the relative binning (RB) likelihood) capable of handling the same within the framework of Bilby. We validate the approximate likelihoods by comparing them with the exact likelihood for a GW170817-like signal (signal-to-noise ratio   1900) containing higher-order modes of radiation. We use the relative-binning likelihood to perform parameter estimation (PE) for a GW170817-like signal, including Earth-rotation effects, detector-size effects, and higher-order modes. We study the system in several detector networks consisting of a single 40 km Cosmic Explorer, a 20 km CE and a present-generation detector at A+ sensitivity. The PE runs with RB take around a day to complete on a typical cluster.
@@ 1 Introduction
The discovery of gravitational waves (GWs) has enabled the study of compact binary coalescences (CBCs) comprising black holes and neutron stars [1, 2, 3, 4, 5]. The LIGO-Virgo-KAGRA (LVK) [6, 7, 8] collaboration has confidently detected around 90 compact CBCs with 200 more candidates from the present observing run [9]. These detections have tested our understanding of gravity, cosmology, and astrophysics [10, 11, 12]. The proposed next-generation ground-based detectors, such as the Cosmic Explorer (CE) [13] and Einstein Telescope (ET) [14], promise an exciting era of gravitational wave detections, offering sensitivities that extend well beyond current capabilities. The next-generation observatories are sensitive enough to detect signals from sources at cosmological distances and provide precise measurements of their parameters which are expected to play a pivotal role in advancing our understanding of fundamental physics, astrophysics, and cosmology [15, 16, 17].
@@
To understand the full potential of these next-generation detectors we need estimates of how well various parameters can be measured. Using conventional Bayesian parameter estimation (PE) is challenging, given the large number of detections ($\mathcal{O}\sim 10^{5}-10^{6}$) [18] per year, some with signal-to-noise ratios (SNRs) greater than 1000. Moreover, signals are long and we need to incorporate corrections due to Earth’s rotation at low frequencies and detector-size at high frequencies [19, 20]. Ignoring these effects is known to cause biases at high SNRs [21]. Higher-order modes of radiation become non-negligible at higher SNRs making Bayesian PE more complex and computationally expensive.
@@
Fisher information matrices offer a computationally cheap way to estimate the uncertainties in measured parameters. Effects on localization due to Earth’s rotation have been extensively studied in this framework [22, 23]. Several constraints on cosmological parameters have also been given using Fisher matrices [24]. This method has also been used for parameter estimation of intermediate mass black holes in a network of detectors and the results agree with Bayesian parameter estimation to an order of magnitude in the parameter space of interest [25]. However, no general Fisher matrix approach taking into account all the effects due to Earth’s rotation, the detector-size and higher modes exist in literature. Also, this approximation becomes non-trivial if the parameter space has multimodalities.
@@
Bayesian parameter estimation does not suffer from the shortcomings of Fisher information matrices but is computationally expensive. A recent paper by Baker et. al. [26] highlights the computational challenges of inference in next-generation detectors. Nearly all Bayesian PE in next-generation detectors involves approximating the likelihood. Pre-merger localization of compact-binary sources in next-generation networks [27] have been studied using relative binning [28]. Reduced order models taking into account only the amplitude modulations due to Earth-rotation using BNS signals lasting 90 minutes in-band from 5 Hz to 2048 Hz have been constructed for a network of detectors [29]. A recent paper highlights challenges Bayesian PE with effects due to Earth’s rotation and detector-size effects have been implemented [21] in the framework of the multibanding approximation [30]. Bayesian PE with relative binning has also been used to measure tidal deformabilities in ET [31]. Similar work has been done for Laser Interferometer Space Antenna (LISA) [32].
@@
In this paper, we develop a generalized framework using Bilby [33] to generate waveforms including effects due to Earth-rotation, finite-size of detectors and higher modes. We develop four new likelihood classes: the exact likelihood including all effects; the multibanding likelihood capable of handling all effects; a relative binning likelihood capable of handling the dominant mode of radiation; a mode-by-mode relative binning likelihood capable of handling all modes of radiation. We do illustrative PE for two GW170817-like [3] sources in four network configurations: a single 40 km CE, a network of LIGO-India and a 40 km CE; a network of a 40 km long and a 20 km long CE; and a network comprising of all detectors. We study the roles played by the higher-order multiples of radiation in these networks.
The rest of the paper is organized as follows. In section 2, we describe our implementations and provide code snippets describing how to use them. In the next section 3 we describe the sources, waveform family and the detection networks used in the study. Finally we validate our methods and present the results of our study in section 4.
@@ 2.1 Frequency-dependent antenna responses
The GW polarization tensors ($e^{+}_{ij},e^{\times}_{ij}$) can be expressed in terms of basis vectors in the transverse plane ($e^{1}_{i},e^{2}_{j}$) given by,
@@
The detected GW strain as a function of time ($t$) can be expressed as $d(t)=h_{D}(t)+n(t)$, where $n$ represents noise and $h_{D}$ is the GW signal at the detector obtained by the product of the detector tensor ($D^{ij}$) [35] and the metric perturbations. In terms of the polarization components, the strain at the detector is expressed as,
$h_{D}(t-t_{c};\Theta)=\sum_{i=+,\times}F_{i}(t-t_{c},f;\xi,I)h_{i}(t-t_{c};\theta,D_{L},\theta_{JN},\phi_{c})$ (4)
@@
$F_{+}(t-t_{c},f;\xi,I):=D^{ij}(t-t_{c},f;\xi,I)e^{+}_{ij}(t-t_{c};\xi)$ (5)
$F_{\times}(t-t_{c},f;\xi,I):=D^{ij}(t-t_{c},f;\xi,I)e^{\times}_{ij}(t-t_{c};\xi)$ (6)
Here $f$ denotes the frequency, $t_{c}$ is the arrival time of the frequency corresponding to the merger at the geocenter and $I$ is the position and orientation of the interferometer. The angle uniquely defining $e^{1}_{i},e^{2}_{j}$ in the transverse plane is the polarization angle ($\psi$). The right ascension ($\rm RA$), declination ($\rm dec$), $\psi$ and $t_{c}$ is collectively denoted by $\xi:=({\rm RA},{\rm dec},\psi,t_{c})$.
@@
The signals are short-lived for current-generation ground-based detectors and so $t\simeq t_{c}$, where $t_{c}$ is the arrival time of the GW at the detector. The detector tensor is also a constant tensor with no dependence on frequency or the direction of the source ($\xi$). Thus for the present generation detectors,
$F_{+}(0,0;\xi,I):=D^{ij}(0,0;I)e^{+}_{ij}(0;\xi)$ (7)
$F_{\times}(0,0;\xi,I):=D^{ij}(0,0;I)e^{\times}_{ij}(0;\xi)$ (8)
@@
Thus, for a given pair of source and a detector the antenna-response functions $F_{+}$ and $F_{\times}$ are constants with no frequency dependence.
Loud sources in next-generation detectors observed from a frame co-rotating with the Earth (the rest-frame of ground-based detectors) shift in the sky, changing the antenna-response function as the signal evolves. The waveform polarizations depend on the intrinsic source parameters like chirp mass ($\mathcal{M}$), mass ratio ($q$) and aligned-spin parameters ($\chi_{1},\chi_{2}$) collectively denoted by $\theta:={\mathcal{M},q,\chi_{1}\chi_{2}}$, the luminosity distance $D_{L}$, the inclination angle $\theta_{\rm JN}$ and the coalescence phase $\phi_{c}$. Collectively, these parameters are represented as $\Theta:=(\xi,\theta,D_{L},\theta_{\rm JN},\phi_{c})$. The earth’s rotation also introduces an integrated Doppler shift as the time delay between the detector and the geocenter is different for various frequencies corresponding to the CBC signal since lower frequencies arrive at earlier times. The effect on the beam patterns and the Doppler effect combined are called Earth-rotation effects in this paper.
@@
The long-wavelength approximation breaks down for a 40 km long detector at high frequencies. For finite wavelengths, the measured strain depends not only on when wavefronts pass through the vertex of the interferometer, but also when it passes through the ends of the arms [19, 20], which introduces an additional dependence on the direction of the wave propagation which in turn depends on the (time-changing) wavelength (or frequency).This effect is referred to as the detector-size effect in this paper.
@@
Thus, for current-generation detectors we need the complete time-frequency dependence as in equation 12, with the $D^{ij}$ having additional frequency and directional dependence, $e^{+,\times}_{ij}$ having temporal dependence and an integrated Doppler shift. Ignoring these effects may introduce systematic biases in parameter estimation (PE) [21].
@@
Time ($\tau:=t-t_{c}$) and frequency are not independent parameters for a GW from a CBC and are related by an invertible function. Parameter estimation is often performed in the frequency domain, necessitating the Fourier transformation of equation 12 to the frequency domain. Analytical computation of the Fourier transform is challenging without approximations. We assume that $h_{i}$’s oscillate rapidly compared to the antenna responses (the stationary phase approximation). We can therefore, take the Fourier transform of the waveform polarizations and convert the times in the antenna response to frequency using the time-frequency relation for a CBC signal computed up to second post-Newtonian (PN) [36] order. While this method is approximate, it provides sufficient accuracy for practical applications. Details of this method can be found in [21].
@@
$h_{D}(f;\Theta)=\sum_{i=+,\times}\sum_{m=-l}^{l}F_{i}(\tau(m,f),f;\xi)\sum_{l=2}^{\infty}h^{lm}_{+}(f;\theta,D_{L},\theta_{\rm JN},\phi_{c},t_{c})$ (9)
where the real-valued $h^{lm}_{+}$ and $h^{lm}_{\times}$ are related to the complex valued $h^{lm}_{+}(f)$ and the spin-2 weighted spherical harmonics ${}_{-2}Y_{lm}$ as
$h^{lm}_{+}(f;\theta,D_{L},\theta_{\rm JN},\phi_{c},t_{c})-ih^{lm}_{\times}(f;\theta,D_{L},\theta_{\rm JN},\phi_{c},t_{c})=_{-2}Y_{lm}(\theta_{\rm JN},\phi_{c})h_{lm}(f;\theta,D_{L},t_{c})$ (10)
In PN theory the time-to-merger for an azimuthal mode $m$ from a frequency $f$, is equivalent to the time to merger of a $m=2$ mode from a frequency $2f/m$ [37, 38]. So the antenna-responses when expressed in terms of frequencies depend on the azimuthal wave number $m$. We implement all functions and classes this paper requires in a cloned version of Bilby. The time-frequency relation for mode $m$ is implemented by a function named calculate_time_to_merger_for_any_mode in gw.utils and this is typically 30 times faster than the method calculate_time_to_merger tested on an Apple M3 Pro. The antenna patterns can be quickly computed as follows:
1$>$ times_to_coalescence = gwutils.calculate_time_to_merger_for_any_mode(frequencies, mass1, mass2, chi1, chi2, mode=2, safety=1)
2$>$ fps, fcs = ifo.frequency_dependent_antenna_response(ra, dec, time, psi, times_to_coalescence, frequencies, start_time, earth_rotation_time_delay = True, earth_rotation_beam_patterns = True, finite_size =True)
The arrays fps and fcs represent the plus and cross-polarization components, corresponding to the frequencies in the frequencies array. Parameters mass1, mass2, chi1, chi2,RA, dec, time, psi, and start_time denote the primary mass, secondary mass, primary spin, secondary spin, right ascension, declination, geocentric time, polarization angle, and the signal start time, respectively. The start time is the difference between the geocentric time and the duration of the signal plus any user-defined offset. The array times_to_coalescence specifies the times to merger for each frequency. The remaining parameters are boolean values set to true if the user wants all effects in the antenna responses. The waveform can be generated as follows:
1$>$ waveform_generator = bilby.gw.WaveformGenerator(duration, sampling_frequency, frequency_domain_source_model=bilby.gw.source.lal_binary_black_hole_individual_modes, parameter_conversion=bilby.gw.conversion.convert_to_lal_binary_black_hole_parameters, waveform_arguments=dict(waveform_approximant=waveformname, reference_frequency=reference_frequency, minimum_frequency=minimum_frequency, mode_array=mode_array))
2$>$ ifo = bilby.gw.detector.InterferometerList([‘CE’])[0]
3$>$ frequencies = waveform_generator.frequency_array
4$>$ parameters, _ = waveform_generator.parameter_conversion(injection_parameters)
5$>$ waveform_polarizations = waveform_generator.frequency_domain_strain(converted_injection_parameters)
6$>$ h[idxs_above_minimum_frequency] = ifo.get_detector_response_for_frequency_dependent_antenna_response(waveform_polarizations, parameters, start_time, frequencies, earth_rotation_time_delay, earth_rotation_beam_patterns,finite_size)’
Here bilby.gw.source.lal_binary_black_hole_individual_mode is a new frequency domain source model that creates a dictionary of the plus and cross polarizations for every mode of radiation. For computational efficiency, modes with the same azimuthal number $m$ are grouped, as they share the same antenna response. An exception is made for relative binning, where all modes are processed individually, as discussed in later sections.
2.2 Sampling the Likelihood

### v2

# Parameter estimation of gravitational-wave signals with frequency-dependent antenna responses and higher modes

@@ Abstract
We implement frequency-dependent antenna responses and develop likelihood classes (standard likelihood, multibanded likelihood, and the relative binning (RB) likelihood) capable of handling the same within the framework of Bilby. We validate the approximate likelihoods by comparing them with the exact likelihood for a GW170817-like signal (luminosity distance scaled such that signal-to-noise ratio $\sim$ 1000) containing higher-order modes of radiation. We use the RB likelihood to perform parameter estimation (PE) for a GW170817-like signal, including Earth-rotation effects, detector-size effects, and higher-order modes. To the best of our knowledge, we are the first to perform proof-of-concept Bayesian parameter estimation, including all the above-mentioned effects. We study the system in several detector networks consisting of a single 40 km Cosmic Explorer, a 20 km CE, and a present-generation detector at A+ sensitivity. The PE runs with RB take around a day to complete on a typical cluster. We find that higher-order modes significantly improve sky localization in scenarios where only a single next-generation detector is operating. This advantage diminishes when a second detector is added to the network; however, in such cases, we observe improved recovery of the polarization angle. While ignoring higher-order modes leads to a reduction in precision, it does not introduce bias in parameter recovery. In contrast, neglecting the effects of Earth’s rotation and the finite size of long-arm detectors can result in biased parameter estimates.
@@ 1 Introduction
The discovery of gravitational waves (GWs) has enabled the study of compact binary coalescences (CBCs) comprising black holes and neutron stars [1, 2, 3, 4, 5]. The LIGO-Virgo-KAGRA (LVK) [6, 7, 8] collaboration has confidently detected around 90 compact CBCs with 250 more candidates from the present observing run [9]. These detections have tested our understanding of gravity, cosmology, and astrophysics [10, 11, 12]. The proposed next-generation ground-based detectors, such as the Cosmic Explorer (CE) [13] and Einstein Telescope (ET) [14], promise an exciting era of gravitational wave detections, offering sensitivities that extend well beyond current capabilities. The next-generation observatories are sensitive enough to detect signals from sources at cosmological distances and provide precise measurements of their parameters, playing a pivotal role in advancing our understanding of fundamental physics, astrophysics, and cosmology [15, 16, 17].
@@
To fully understand the potential of these next-generation detectors, we require estimates of how accurately various parameters can be measured. Using conventional Bayesian parameter estimation (PE) is challenging, given the large number of detections ($\mathcal{O}\sim 10^{5}-10^{6}$) [18] per year, some with signal-to-noise ratios (SNRs) greater than 1000. Moreover, signals are long and we need to incorporate corrections due to Earth’s rotation at low frequencies and detector-size at high frequencies [19, 20]. Ignoring these effects is known to cause biases at high SNRs [21]. Higher-order modes of radiation become non-negligible at higher SNRs, making Bayesian PE more complex and computationally expensive.
@@
Fisher information matrices offer a computationally cheap way to estimate the uncertainties in measured parameters. Effects on localization due to Earth’s rotation have been extensively studied in this framework [22, 23]. Several constraints on cosmological parameters have also been given using Fisher matrices [24]. This method has also been used for parameter estimation of intermediate mass black holes in a network of detectors and the results agree with Bayesian parameter estimation to an order of magnitude in the parameter space of interest [25]. However, no general Fisher matrix approach taking into account all the effects due to Earth’s rotation, the detector size, and higher modes exists in the literature. Also, this approximation becomes invalid if the parameter space has multimodalities.
@@
Bayesian parameter estimation does not suffer from the shortcomings of Fisher information matrices but is computationally expensive. A recent paper by Baker et al. [26] highlights the computational challenges of inference in next-generation detectors. Nearly all Bayesian PE in next-generation detectors involves approximating the likelihood due to computational constraints [27]. Pre-merger localization of compact-binary sources in next-generation networks [27] has been studied using relative binning [29, 28]. Reduced-order models that take into account only the amplitude modulations due to Earth rotation using BNS signals lasting 90 minutes in-band from 5 Hz to 2048 Hz have been constructed for a network of detectors [29]. Bayesian PE with Earth’s rotation effects and detector-size effects for long BNS signals have been implemented by Baral et al. [21] in the framework of the multibanding approximation [30]. Bayesian PE with relative binning has also been used to measure tidal deformabilities in ET [31]. Similar work has been done for Laser Interferometer Space Antenna (LISA) [32].
@@
In this paper, we develop a generalized framework using Bilby [33] to generate waveforms that include effects due to Earth rotation, the finite size of detectors, and higher modes. We develop four new likelihood classes: the exact likelihood including all effects; the multibanding likelihood capable of handling all effects; a relative binning likelihood capable of handling the dominant mode of radiation; a mode-by-mode relative binning likelihood capable of handling all modes of radiation. We do illustrative PE for two GW170817-like [3] sources in four network configurations: a single 40 km CE, a network of LIGO-India and a 40 km CE; a network of a 40 km long and a 20 km long CE; and a network comprising all detectors. The luminosity distance has been scaled to fix the SNR to $\sim$ 1000. the We study the roles played by the higher-order multiples of radiation in these networks.
The rest of the paper is organized as follows. Section 2 describes the methods implemented for this study. Section 3 details the gravitational wave sources, waveform families, and detection networks used in our analysis. Finally, Section 4 presents the validation of our methods and the main results of the study.
@@ 2.1 Frequency-dependent antenna responses
The GW polarization tensors ($e^{+}_{ij},e^{\times}_{ij}$) can be expressed in terms of basis vectors in the transverse plane ($e^{a}_{i}$) given by,
@@
Here $i=x,y$ runs over the spatial indices on the transverse plane with the z-axis being the normal vector from the detector to the source, and $a=1,2$ denote the arms of the interferometer. Thus, loud sources in next-generation detectors observed from a frame co-rotating with the Earth (the rest-frame of ground-based detectors) shift in the sky, as the signal evolves. The Earth’s rotation also introduces an integrated Doppler shift. This is because the time delay between the detector and the geocenter is different for various frequencies corresponding to the CBC signal, since lower frequencies arrive at earlier times. The effect on the beam patterns and the Doppler effect combined is called the Earth-rotation effects in this paper.
The detected GW strain as a function of time ($t$) can be expressed as $d(t)=h_{D}(t)+n(t)$, where $n$ represents noise and $h_{D}$ is the GW signal at the detector, obtained by the product of the detector tensor ($D^{ij}$) [35] and the metric perturbations. The detector tensor is given by,
$D_{ij}(t,f_{L};\xi,I):=\frac{1}{2}\Big[D_{1}(t,f_{L};\xi,I)e^{1}_{i}(t;\xi,I)e^{1}_{j}(t;\xi,I)-D_{2}(t,f_{L};\xi,I)e^{2}_{i}(t;\xi,I)e^{2}_{j}(t;\xi,I)\Big]$ (4)
where the detector scalar ($D_{a}$) [19, 20] is given by
$D_{a}(t,f_{L};\xi,I)$ $=\exp\{{\pi if_{L}(1-\textbf{z}\cdot\textbf{e}^{a})}\}\text{sinc}\{\pi f_{L}(1+\textbf{z}\cdot\textbf{e}^{a})\}$ (5)
$+\exp\{{-\pi if_{L}(1+\textbf{z}\cdot\textbf{e}^{a})}\}\text{sinc}\{\pi f_{L}(1-\textbf{z}\cdot\textbf{e}^{a})\}$
Here $f$ denotes the frequency, $I$ is the position and orientation of the interferometer $f_{L}:=fL/c$ and $\text{sinc }x:=\sin x/x$. The angle uniquely defining $e^{a}_{i}$ in the transverse plane is the polarization angle ($\psi$). The right ascension ($\rm RA$), declination ($\rm dec$), $\psi$ is collectively denoted by $\xi:=({\rm RA},{\rm dec},\psi)$.
In terms of the polarization components, the strain at the detector is expressed as,
$h_{D}(t;\Theta)=\sum_{i=+,\times}F_{i}(t,f_{L};\xi,t_{c},I)h_{i}(t-t_{c};\theta,D_{L},\theta_{JN},\phi_{c})$ (6)
@@
$F_{+}(t,f_{L};\xi,I):=D^{ij}(t,f_{L};\xi,I)e^{+}_{ij}(t;\xi,I)$ (7)
$F_{\times}(t,f_{L};\xi,I):=D^{ij}(t,f_{L};\xi,I)e^{\times}_{ij}(t;\xi,I)$ (8)
Here $t_{c}$ is the arrival time of the frequency corresponding to the merger at the geocenter. The waveform polarizations ($h_{+},h_{\times}$) depend on the intrinsic source parameters like chirp mass ($\mathcal{M}$), mass ratio ($q$) and aligned-spin parameters ($\chi_{1},\chi_{2}$), collectively denoted by $\theta:=({\mathcal{M},q,\chi_{1}\chi_{2}})$, the luminosity distance $D_{L}$, the inclination angle $\theta_{\rm JN}$ and the coalescence phase $\phi_{c}$. Collectively, these parameters are represented as $\Theta:=(\xi,\theta,D_{L},\theta_{\rm JN},\phi_{c},t_{c})$.
@@
The signals are short-lived for current-generation ground-based detectors and so $t\simeq t_{c}$, where $t_{c}$ is the arrival time of the GW at the detector. Thus, the polarization tensor does not vary with time. The term $f_{L}\rightarrow 0\implies D_{ij}\rightarrow 1$ (see equation 5) for present detectors as the wavelength of the GW is long compared to the size of the detector. The long-wavelength approximation makes the detector tensor a constant tensor with no dependence on frequency or the direction of the source. Thus, for the present generation detectors,
$F_{+}(t_{c},0;\xi,I):=e^{+}_{ij}(t_{c};\xi,I)$ (9)
$F_{\times}(t_{c},0;\xi,I):=e^{\times}_{ij}(t_{c};\xi,I)$ (10)
@@
Thus, for a given pair of source and detector, the antenna-response functions $F_{+}$ and $F_{\times}$ are constants with no frequency dependence.
@@
The long-wavelength approximation breaks down for a 40 km long detector at high frequencies. For finite wavelengths, the measured strain depends not only on when wavefronts that pass through the vertex of the interferometer, but also when it passes through the end of the arms [19, 20]. This introduces an additional dependence on the direction of the wave propagation, which in turn depends on the (time-changing) wavelength (or frequency). This effect is referred to as the detector-size effect in this paper.
@@
Thus, for next-generation detectors we need the complete time-frequency dependence as in equation 7, with the $D^{ij}$ having additional frequency and directional dependence, $e^{+,\times}_{ij}$ having temporal dependence and an integrated Doppler shift. Ignoring these effects may introduce systematic biases in parameter estimation (PE) [21].
@@
Parameter estimation is often performed in the frequency domain, necessitating the Fourier transformation of equation 14 to the frequency domain. Analytical computation of the Fourier transform is challenging without approximations. We first note that for a particular quasicircular inspiral mode of a GW from a CBC time ($\tau:=t-t_{c}$) and frequency are not independent parameters and are related by an invertible function. We assume that $h_{i}$’s oscillate rapidly compared to the antenna responses (the stationary phase approximation). We can therefore, take the Fourier transform of the waveform polarizations and convert the times in the antenna response to frequency using a frequency to $\tau$ converter ($\tau(m,f)$)  [^1] In PN theory the time-to-merger for an azimuthal mode $m$ from a frequency $f$, is equivalent to the time to merger of a $m=2$ mode from a frequency $2f/m$ [37, 38] correct up to 2 PN for an arbitrary azimuthal mode as described in [36] While this method is approximate, it provides sufficient accuracy for practical applications. Details of this method can be found in [21].
@@
$h_{D}(f;\Theta)=\sum_{i=+,\times}\sum_{|m|=0}^{\infty}F_{i}(t(\tau(m,f),t_{c}),f_{L};\xi,t_{c},I)\sum_{\begin{subarray}{c}{l=2}\\ {l>m}\end{subarray}}^{\infty}\tilde{h}^{lm}_{i}(f;\theta,D_{L},\theta_{\rm JN},\phi_{c},t_{c})$ (11)
where $\tilde{h}^{lm}_{+}$ and $\tilde{h}^{lm}_{\times}$ are the fourier transforms of the real-valued $h^{lm}_{+}$ and $h^{lm}_{\times}$, which are related to the complex valued $h_{lm}(t)$ and the spin-2 weighted spherical harmonics $_{-2}Y_{lm}$ as
$h^{lm}_{+}(t;\theta,D_{L},\theta_{\rm JN},\phi_{c},t_{c})-ih^{lm}_{\times}(t;\theta,D_{L},\theta_{\rm JN},\phi_{c},t_{c})=_{-2}Y_{lm}(\theta_{\rm JN},\phi_{c})h_{lm}(t;\theta,D_{L},t_{c})$ (12)
We implement the classes and functions required to generate a waveform with higher modes and frequency-dependent antenna response in a forked version of Bilby. To summarize, this involves creating a new frequency domain source model that creates a dictionary of the plus and cross polarizations for every mode of radiation, functions capable of generating frequency-dependent antenna responses for any mode. For computational efficiency, modes with the same azimuthal number $m$ are grouped, as they share the same antenna response. An exception is made for relative binning, where all modes are processed individually, as discussed in later sections.
2.2 Sampling the Posterior